OpenStudy (anonymous):

The table shows the heights of four people Karyl plays basketball with. A data set shows various players' heights in inches: sixty-four, sixty-five, sixty-seven, sixty-eight What will happen to the mean height when a player who is 73 inches tall joins the group?

OpenStudy (anonymous):

It will increase because 73 is greater than all the heights.

OpenStudy (anonymous):

Sorry, The answer choices are: The mean increases by one point four inches. The mean increases by eighteen point two five inches. The mean does not change. The mean decreases by thirteen point two inches. What one?

OpenStudy (anonymous):

oh ok. Original mean=(64+65+67+68)/4, which is 66. New mean=(64+65+67+68+73)/5, which is 67.4. The mean increases by 1.4 inches.

OpenStudy (anonymous):

Thank you!